OpenStudy (anonymous):

a recipe calls for 2/5 cup of chicken broth and 1/6 cup of vegetable broth . what is the total amount of broth need for the recipe ? help plz

OpenStudy (rulnick):

2/5 + 1/6 = 12/30 + 5/30 = 17/30 17/30 cup